# Interactive Query Answering on Knowledge Graphs with Soft Entity Constraints

### v3

@@ Abstract
Methods for query answering over incomplete knowledge graphs retrieve entities that are likely to be answers, which is particularly useful when such answers cannot be reached by direct graph traversal due to missing edges. However, existing approaches have focused on queries formalized using first-order-logic. In practice, many real-world queries involve constraints that are inherently vague or context-dependent, such as preferences for attributes or related categories. Addressing this gap, we introduce the problem of query answering with soft constraints. We formalize the problem and introduce two efficient methods designed to adjust query answer scores by incorporating soft constraints without disrupting the original answers to a query. These methods are lightweight, requiring tuning only two parameters or a small neural network trained to capture soft constraints while maintaining the original ranking structure. To evaluate the task, we extend existing QA benchmarks by generating datasets with soft constraints. Our experiments demonstrate that our methods can capture soft constraints while maintaining robust query answering performance and adding very little overhead. With our work, we explore a new and flexible way to interact with graph databases that allows users to specify their preferences by providing examples interactively.
@@ 1 Introduction
Knowledge graphs (KGs) are graph-structured databases that store information about a domain using triples of the form (subject, predicate, object). Such a compact representation enables efficient algorithms for answering queries about entities over the domain encoded by the KG (fensel_knowledge_2020; hogan_knowledge_2021). In principle, answers to these queries can be obtained by traversing the graph and collecting the entities that satisfy the given conditions. However, this strategy misses answers due to connections that are not explicit in a KG that may be incomplete, but that are likely to satisfy a query, for example because of their similarity to known answers. The problem of finding likely answers to simple queries over KGs has been studied in areas such as rule mining (galarraga_amie_2013; meilicke_fine-grained_2018; meilicke_anytime_2019; qu_rnnlogic_2020) and link prediction (nickel_review_2015; ji_survey_2020), and complex query answering (hamilton2018gqe; daza2020message; ren_query2box_2020; ren_beta_2020; arakelyan2021cqd; arakelyan2023cqdA; bai2023qto; zhu_neural-symbolic_2022; ren2024neural). Given a query, these methods score all the entities in the KG, indicating the likelihood of meeting all conditions specified in the query. Notably, prior works have focused on conditions that can be specified using different subsets of first order logic.
@@ 2 Related Work
We build on methods for learning on KGs that retrieve query answers that are not reachable via graph traversal. Early work focused on simple link prediction, where an edge is predicted between two entities (nickel2011rescal; bordes2013transe; trouillon2016complex; lacroix2018complex_n3; sun2019rotate). More recently, the set of queries that learning-based methods can tackle has been broadened to complex queries (e.g., paths of length two or three with intermediate variables), using a formulation of queries based on a small (hamilton2018gqe; daza2020message; ren_query2box_2020; ren_beta_2020; arakelyan2021cqd; arakelyan2023cqdA; bai2023qto; zhu_neural-symbolic_2022; ren2024neural) or a larger  (cucumides_unravl_2024; yin_rethinking_2024) subset of first order logic. In all these cases, the output of a query answering model is a score for each entity in the graph that indicates how likely it is to meet constraints specified using first order logic. Our work complements these approaches by supporting soft constraints that cannot be expressed in the same way, due to the inherent limitations of the language of first order logic.
@@
Active Learning on KGs. The idea of collecting entities that exemplify soft constraints is closely related to the problem of active learning (fu_survey_2013; ren_survey_2022). Active learning allows training models incrementally when data is of limited availability. In the context of KGs, it has been applied to sampling training data for link prediction models under a constrained computational budget (ostapuk_activelink_2019), error detection (dong_active_2023), and KG alignment (huang_deep_2023); all cases in which a prediction is well-defined. Applying active learning directly to an existing query answering model in order to capture soft constraints is not trivial, as a single query can have different types of soft constraints (as illustrated in Fig. 1), potentially leading to contradicting model updates. We avoid this by training a lightweight neural query reranker via supervised learning, using automatically generated data of soft preferences.
@@
Information Retrieval (IR) and Recommender Systems. A common task in IR is to retrieve documents relevant to a given query (manning2009introduction), for example based on their similarity to a textual query using sequence embeddings (kenton2019bert). We adopt a similar approach in one of the baselines in our experiments. We also draw inspiration from Learning to Rank (LTR), where models are trained to order items according to their relative relevance. Pairwise methods such as RankNet and LambdaRank (burges_learning_2005; burges_learning_2006) and their boosted extension LambdaMART (burges2010ranknet) learn from preference pairs by directly optimizing ranking metrics such as NDCG, and have been shown to outperform regression- or classification-based ranking approaches. Modern implementations, such as the one provided in LightGBM (ke2017lightgbm), efficiently extend these methods to large datasets while supporting graded relevance labels and fast tree-based inference. These approaches generate a new ranking from scratch given a set of preferences; we, on the other hand, aim to preserve the original score relationships while biasing the ranking towards soft constraints. Other learning-to-rank methods take a list-wise approach that considers permutations of candidate lists (cao_learning_2007; xia_listwise_2008; zhu_listwise_2020), but these approaches do not scale well to the number of entities typically found in a knowledge graph. Lastly, the field of Recommender Systems has explored the use of interaction data (e.g., clicks, ratings, and browsing history) to generate recommendations (koren2021advances). Various collaborative filtering models exist in the literature (ning2011slim; steck2021negative; koren2009matrix; rendle2012bpr; he2017neural; christoffel2015blockbusters), some of which demonstrate the effectiveness of neural networks at capturing preference data.
@@ 3 Preliminaries
Answering this query requires assigning entities from $\mathcal{V}$ to variables $v_{1}$ and $v_{2}$ such that the conjunction specifying the constraints is true according to the information contained in the KG. More generally, and following the notation in bai2023qto, we consider first-order logic (FOL) queries that can include conjunctions, disjunctions, and negations, which we write in disjunctive normal form as follows:
@@
Approximate Query Answering. In order to provide answers to queries that are not limited to the information explicitly stated by the edges in the graph, recent works have introduced machine learning models for approximate query answering (QA) (hamilton2018gqe; daza2020message; ren_query2box_2020; ren_beta_2020; arakelyan2021cqd; arakelyan2023cqdA; bai2023qto; zhu_neural-symbolic_2022; ren2024neural; cucumides_unravl_2024; yin_rethinking_2024). For a given query, these methods compute a score for all the entities in the KG according to how likely they are to be answers to a query. While the specific mechanisms for achieving this vary among methods, in general we denote a QA model as a function $a$ that maps a query $q$ to a vector $a(q)\in\mathbb{R}^{|\mathcal{V}|}$ of real-valued scores, with one score for each entity in the KG.
@@ 4 Query Answering with Soft Constraints
Definition 1 (Soft constraint (informal)).
A soft constraint is a loose requirement based on semantic similarity, which can guide answers to a query towards desirable types of answers without imposing strict logical criteria. For example, the soft constraint “awards involving performers” suggests prioritizing answers related to acting awards, without strictly excluding other awards.
We consider a scenario where soft constraints are specified by means of specific examples of entities that represent the desired (or undesired) type of answer. We formalize this through labeled preferences. A preference is a tuple $(e,l)$, where $e\in\mathcal{V}$ is an entity in the KG, and $l\in\{0,1\}$ is a label indicating whether entities similar to $e$ should be preferred ($l=1$) or de-emphasized ($l=0$).
Since specifying an exhaustive list of preferences can be impractical, we consider an interactive setting in which preferences are provided incrementally.
@@
Let $q$ be a query with hard logical constraints evaluated over a knowledge graph, and let $a^{(0)}(q)$ denote the initial list of scores produced by a query answering model. In the interactive query answering with soft constraints setting, the scores are adjusted iteratively based on a growing set of user-provided preferences: after $t$ interactions, the preference set is
$P(t)=\{(e_{1},l_{1}),\ldots,(e_{t},l_{t})\}.$ (3)
At each iteration $t\geq 1$, we denote the adjusted list of scores as
$a^{(t)}(q)=f\big(a^{(0)}(q),P(t)\big),$ (4)
where $f$ is a function that updates the initial scores according to the accumulated preferences.
@@ 5 Incorporating Soft Constraints
The majority of existing QA models learn entity embeddings for computing query scores (hamilton2018gqe; daza2020message; ren_query2box_2020; ren_beta_2020; arakelyan2021cqd; arakelyan2023cqdA; bai2023qto; zhu_neural-symbolic_2022, among others). We propose to reuse the embeddings of the base QA model (which are already required to compute the initial vector of scores $a(q)$) to compute the cosine similarity $\operatorname{sim}(e_{i},e)$ required in Equation˜9.
@@
We now introduce two implementations of the update function $f$ in Equation˜8: a linear combination with fixed coefficients, and a neural reranker where the coefficients are computed conditionally.
@@
Motivated by prior work on reranking methods for search over knowledge graphs (gerritse2020graph; daza2021blp), we first propose a score update that is computed as a convex combination of the original score, and a linear combination of the average cosine similarities with $P^{+}$ and $P^{-}$. We refer to this as the Cosine update:
@@
The cosine update thus requires tuning two parameters, $\alpha\in(0,1)$ and $\beta\in(-1,1)$ that control the weight given to the original score, and the average cosine similarities, which can be done using a validation set.
@@
We can further adapt the weights given to each term in the score update function in Equation˜10 by computing them with a small multi-layer perceptron (MLP) with four inputs: the original score $a[e]$, the average similarities $\delta(P^{+},e)$ and $\delta(P^{-},e)$, and the embedding $\mathbf{e}\in\mathbb{R}^{d}$ of entity $e$ (which we reuse from the base QA model and do not fine-tune).
Intuitively, making the weights conditional allows to contextualize the score update to specific values of scores, similarities, and the entity as captured by its embedding. The MLP outputs two scalars $\alpha_{p},\alpha_{n}\in(0,1)$, which we then use in the update rule. We refer to this as the Neural Query Reranker (NQR) update:
@@
To achieve this, we adopt the RankNet loss (burges_learning_2005), which encourages pairs of relevant and non-relevant items to be ordered correctly. Given two entities $e_{i}$ and $e_{j}$ with scores $a[e_{i}]$ and $a[e_{j}]$, RankNet minimizes the binary cross-entropy over the probability $\sigma(a[e_{i}]-a[e_{j}])$ that $e_{i}$ should be ranked above $e_{j}$. For NQR, we define two complementary loss terms:
@@ 6.1 Datasets
Prior work on query answering on KGs has relied on datasets containing query-answer pairs $(q,\mathcal{A})$, where $q$ is a query and $\mathcal{A}\subset\mathcal{V}$ is the set of answers computed over a KG (ren_query2box_2020; ren_beta_2020). We extend these datasets by deriving preference data from observable clusters of entities in the set $\mathcal{A}$. Motivated by prior work showing that textual embeddings correlate with human notions of similarity and relatedness (2019reimerssbert; grand2022semantic; merk2022seeing), we achieve this by clustering embeddings of answers to a query, which we use as a computational proxy of similarity that allows us to study the problem at a scale of thousands of queries and comprising different notions of similarity. Concretely, this process consists of three steps (we provide additional details in Appendix  B):
@@
Figure 2: Example dendrogram used to generate preference sets. Highlighted in green are preferred entities, other entities are non-preferred.
@@
Prior work has relied on subsets of encyclopedic KGs such as NELL (carlson_toward_2010) and Freebase (bollacker_freebase_2008) to benchmark methods for query answering. These KGs contain statements about people, organizations, and locations around the world. For our experiments, we select FB15k237 (toutanova_observed_2015) and the complex queries generated for it by ren_beta_2020. To further explore the potential of query answering in a different domain we use Hetionet (himmelstein_systematic_2017), a biomedical KG containing $\sim$2M edges between genes, diseases, and drugs, among other types. For Hetionet, we extract complex queries of different types, and then for both datasets we run the three steps above to collect preference data. We consider 14 types of first order logic queries of varying complexity, including negation. We present statistics of the final datasets and splits in Table 1. For training NQR, we only use 1-hop, link prediction queries. For the validation and test sets we use the full set of 14 types of complex queries. More detailed statistics can be found in Appendix  C.

### v4

@@ Abstract
Methods for query answering over incomplete knowledge graphs retrieve entities that are likely to be answers, which is particularly useful when such answers cannot be reached by direct graph traversal due to missing edges. However, existing approaches have focused on queries formalized using first-order-logic. In practice, many real-world queries involve constraints that are inherently vague or context-dependent, such as preferences for attributes or related categories. Addressing this gap, we introduce the problem of query answering with soft constraints. We formalize the problem and introduce two efficient methods designed to adjust query answer scores by incorporating soft constraints without disrupting the original answers to a query. These methods are lightweight, requiring tuning only two hyperparameters or a small neural network trained to capture soft constraints while maintaining the original ranking structure. To evaluate the task, we extend existing QA benchmarks by generating datasets with soft constraints. Our experiments demonstrate that our methods can capture soft constraints while maintaining robust query answering performance and adding very little overhead. With our work, we explore a new and flexible way to interact with graph databases that allows users to specify their preferences by providing examples interactively.
Our code is available at https://github.com/dfdazac/nqr.
@@ 1 Introduction
Knowledge graphs (KGs) are graph-structured databases that store information about a domain using triples of the form (subject, predicate, object). Such a compact representation enables efficient algorithms for answering queries about entities over the domain encoded by the KG (Fensel et al., 2020; Hogan et al., 2021). In principle, answers to these queries can be obtained by traversing the graph and collecting the entities that satisfy the given conditions. However, this strategy misses answers due to connections that are not explicit in a KG that may be incomplete, but that are likely to satisfy a query, for example because of their similarity to known answers. The problem of finding likely answers to simple queries over KGs has been studied in areas such as rule mining (Galárraga et al., 2013; Meilicke et al., 2018; 2019; Qu et al., 2020) and link prediction (Nickel et al., 2016; Ji et al., 2022), and complex query answering (Hamilton et al., 2018; Daza and Cochez, 2020; Ren et al., 2020; Ren and Leskovec, 2020; Arakelyan et al., 2021; 2023; Bai et al., 2023; Zhu et al., 2022; Ren et al., 2024). Given a query, these methods score all the entities in the KG, indicating the likelihood of meeting all conditions specified in the query. Notably, prior works have focused on conditions that can be specified using different subsets of first order logic.
@@ 2 Related Work
Recent works have introduced methods for inferring plausible answers that are not reachable via explicit graph traversal, typically due to incompleteness in the graph. We build on methods for learning on KGs that retrieve query answers that are not reachable via graph traversal. Early work focused on simple link prediction, where an edge is predicted between two entities (Nickel et al., 2011; Bordes et al., 2013; Trouillon et al., 2016; Lacroix et al., 2018; Sun et al., 2019). More recently, the set of queries that learning-based methods can tackle has been broadened to complex queries (e.g., paths of length two or three with intermediate variables), using a formulation of queries based on a small (Hamilton et al., 2018; Daza and Cochez, 2020; Ren et al., 2020; Ren and Leskovec, 2020; Arakelyan et al., 2021; 2023; Bai et al., 2023; Zhu et al., 2022; Ren et al., 2024) or a larger  (Cucumides et al., 2024; Yin et al., 2024) subset of first order logic. In all these cases, the output of a query answering model is a score for each entity in the graph that indicates how likely it is to meet constraints specified using first order logic. Our work complements these approaches by supporting soft constraints that cannot be expressed in the same way, due to the inherent limitations of the language of first order logic.
@@
Active Learning on KGs. The idea of collecting entities that exemplify soft constraints is closely related to the problem of active learning (Fu et al., 2013; Ren et al., 2022). Active learning allows training models incrementally when data is of limited availability. In the context of KGs, it has been applied to sampling training data for link prediction models under a constrained computational budget (Ostapuk et al., 2019), error detection (Dong et al., 2023), and KG alignment (Huang et al., 2023); all cases in which a prediction is well-defined. Applying active learning directly to an existing query answering model in order to capture soft constraints is not trivial, as a single query can have different types of soft constraints (as illustrated in Fig. 1), potentially leading to contradicting model updates. We avoid this by training a lightweight neural query reranker via supervised learning, using automatically generated data of soft preferences.
@@
Information Retrieval (IR) and Recommender Systems. A common task in IR is to retrieve documents relevant to a given query (Manning et al., 2008), for example based on their similarity to a textual query using sequence embeddings (Devlin et al., 2019). We adopt a similar approach in one of the baselines in our experiments. We also draw inspiration from Learning to Rank (LTR), where models are trained to order items according to their relative relevance. Pairwise methods such as RankNet and LambdaRank (Burges et al., 2005; 2006) and their boosted extension LambdaMART (Burges, 2010) learn from preference pairs by directly optimizing ranking metrics such as NDCG, and have been shown to outperform regression- or classification-based ranking approaches. Modern implementations, such as the one provided in LightGBM (Ke et al., 2017), efficiently extend these methods to large datasets while supporting graded relevance labels and fast tree-based inference. These approaches generate a new ranking from scratch given a set of preferences; we, on the other hand, aim to preserve the original score relationships while biasing the ranking towards soft constraints.
Other learning-to-rank methods take a list-wise approach that considers permutations of candidate lists (Cao et al., 2007; Xia et al., 2008; Zhu and Klabjan, 2020), but these approaches do not scale well to the number of entities typically found in a knowledge graph.
Lastly, the field of Recommender Systems has explored the use of interaction data (e.g., clicks, ratings, and browsing history) to generate recommendations (Koren et al., 2021). Various collaborative filtering models exist in the literature (Ning and Karypis, 2011; Steck and Liang, 2021; Koren et al., 2009; Rendle et al., 2012; He et al., 2017; Christoffel et al., 2015), some of which demonstrate the effectiveness of neural networks at capturing preference data. Other approaches, such as Rocchio’s algorithm (Rocchio Jr, 1971; Manning et al., 2008), introduce feedback into queries by updating a query vector with a new one as examples are added. However, our problem differs from this setting as the starting point consists of the scores of a base QA model for the unconstrained query, for which no query vectors are available that can be updated.
@@ 3 Preliminaries
Answering this query requires assigning entities from $\mathcal{V}$ to variables $v_{1}$ and $v_{2}$ such that the conjunction specifying the constraints is true according to the information contained in the KG. More generally, and following the notation in Bai et al. (2023), we consider first-order logic (FOL) queries that can include conjunctions, disjunctions, and negations, which we write in disjunctive normal form as follows:
@@
Approximate Query Answering. In order to provide answers to queries that are not limited to the information explicitly stated by the edges in the graph, recent works have introduced machine learning models for approximate query answering (QA) (Hamilton et al., 2018; Daza and Cochez, 2020; Ren et al., 2020; Ren and Leskovec, 2020; Arakelyan et al., 2021; 2023; Bai et al., 2023; Zhu et al., 2022; Ren et al., 2024; Cucumides et al., 2024; Yin et al., 2024). For a given query, these methods compute a score for all the entities in the KG according to how likely they are to be answers to a query. While the specific mechanisms for achieving this vary among methods, in general we denote a QA model as a function $a$ that maps a query $q$ to a vector $a(q)\in\mathbb{R}^{|\mathcal{V}|}$ of real-valued scores, with one score for each entity in the KG.
@@ 4 Query Answering with Soft Constraints
Definition 1 (Soft constraint).
A soft constraint is a boolean predicate $s:\mathcal{V}\rightarrow\{0,1\}$ that indicates whether an entity satisfies a desired semantic property. Furthermore, the predicate cannot be expressed using the logical form of Equation˜2 and is therefore not directly evaluable from the triples in $\mathcal{E}$.
Soft constraints can be used to further restrict the answer set of a query by requiring that the entity assigned to the target variable satisfies the constraint. Let $s:\mathcal{V}\rightarrow\{0,1\}$ be a soft constraint predicate. The constrained answer set of query $q$ is defined as $\hat{A}_{q}=\{e\in\mathcal{A}_{q}\mid s(e)\}$.
Preference sets.
Since the predicate $s$ is not present in the KG and it is not explicitly known, we propose to rely on a small set of labeled examples that can be provided interactively at different time steps $t$. Formally, a preference set is denoted as $P(t)=\{(e_{1},s(e_{1})),\ldots,(e_{k},s(e_{k}))\}$, with $e_{i}\in\mathcal{V}$.
@@
Let $q$ be a query with hard logical constraints evaluated over a knowledge graph, $a^{(0)}(q)$ denote the initial list of scores produced by a query answering model, and $P(t)$ a preference set representing a soft constraint $s$. The problem of interactive query answering with soft constraints consists of computing an adjusted list of scores $a^{(t)}(q)$ that incorporates the information contained in $P(t)$.
The adjusted scores should satisfy two desirable properties.
Preserving of logical constraints.
Entities that do not satisfy the logical query should not be ranked above valid answers: $a^{(t)}(q)[u]>a^{(t)}(q)[v]\quad\forall u\in\mathcal{A}_{q},\;\forall v\notin\mathcal{A}_{q}.$
Capturing preferences.
Among entities satisfying the hard constraints, those that also satisfy the soft constraint should be ranked higher: $a^{(t)}(q)[u]>a^{(t)}(q)[v]\quad\forall u\in\hat{\mathcal{A}}_{q},\;\forall v\in\mathcal{A}_{q}\setminus\hat{\mathcal{A}}_{q}.$
@@ 5 Incorporating Soft Constraints
The majority of existing QA models learn entity embeddings for computing query scores (Hamilton et al. (2018); Daza and Cochez (2020); Ren et al. (2020); Ren and Leskovec (2020); Arakelyan et al. (2021; 2023); Bai et al. (2023); Zhu et al. (2022), among others). We propose to reuse the embeddings of the base QA model (which are already required to compute the initial vector of scores $a(q)$) to compute the cosine similarity $\operatorname{sim}(e_{i},e)$ required in Equation˜7.
@@
We now introduce two implementations of the update function $f$ in Equation˜6: a linear combination with fixed coefficients, and a neural reranker where the coefficients are computed conditionally.
@@
Motivated by prior work on reranking methods for search over knowledge graphs (Gerritse et al., 2020; Daza et al., 2021), we first propose a score update that is computed as a convex combination of the original score, and a linear combination of the average cosine similarities with $P^{+}$ and $P^{-}$. We refer to this as the Cosine update:
@@
Figure 2: To adjust the score of an entity $e\in\mathcal{V}$ (Best Actor), NQR computes similarity scores with entities in the sets $P^{+}$ (Best Actress) and $P^{-}$ (Best Sound) using their embeddings (shown as emb). These scores, together with the original score (here, 0.7) and the embedding of $e$ are the input to an MLP that computes coefficients $\alpha_{p}$ and $\alpha_{n}$ for Equation˜9.
The cosine update thus requires tuning two hyperparameters, $\alpha\in(0,1)$ and $\beta\in(-1,1)$ that control the weight given to the original score, and the average cosine similarities, which can be done using a validation set.
@@
We can further adapt the weights given to each term in the score update function in Equation˜8 by computing them with a small multi-layer perceptron (MLP) with four inputs: the original score $a[e]$, the average similarities $\delta(P^{+},e)$ and $\delta(P^{-},e)$, and the embedding $\mathbf{e}\in\mathbb{R}^{d}$ of entity $e$ (which we reuse from the base QA model and do not fine-tune).
Intuitively, making the weights conditional allows to contextualize the score update to specific values of scores, similarities, and the entity as captured by its embedding. The MLP outputs two scalars $\alpha_{p},\alpha_{n}\in(0,1)$, which we then use in the update rule (see Figure˜2). We refer to this as the Neural Query Reranker (NQR) update:
@@
To achieve this, we adopt the RankNet loss (Burges et al., 2005), which encourages pairs of relevant and non-relevant items to be ordered correctly. Given two entities $e_{i}$ and $e_{j}$ with scores $a[e_{i}]$ and $a[e_{j}]$, RankNet minimizes the binary cross-entropy over the probability $\sigma(a[e_{i}]-a[e_{j}])$ that $e_{i}$ should be ranked above $e_{j}$. For NQR, we define two complementary loss terms:
@@ 6.1 Datasets
Prior work on query answering on KGs has relied on datasets containing query-answer pairs $(q,\mathcal{A})$, where $q$ is a query and $\mathcal{A}\subset\mathcal{V}$ is the set of answers computed over a KG (Ren et al., 2020; Ren and Leskovec, 2020). We extend these datasets by deriving preference data from observable clusters of entities in the set $\mathcal{A}$.
Figure 3: Example dendrogram used to generate preference sets. Highlighted in green are preferred entities.
Motivated by prior work showing that textual embeddings correlate with human notions of similarity and relatedness (Reimers and Gurevych, 2019; Grand et al., 2022; Merkx et al., 2022), we achieve this by clustering embeddings of answers to a query, which we use as a computational proxy of similarity that allows us to study the problem at a scale of thousands of queries and comprising different notions of similarity. Concretely, this process consists of three steps (we provide additional details in Appendix  B):
@@
This results in clusters capturing information not available to the QA model or any of the reranking methods, which rely only on graph embeddings and have no access to textual descriptions.
@@
Prior work has relied on subsets of encyclopedic KGs such as NELL (Carlson et al., 2010) and Freebase (Bollacker et al., 2008) to benchmark methods for query answering. These KGs contain statements about people, organizations, and locations around the world. For our experiments, we select FB15k237 (Toutanova and Chen, 2015) and the complex queries generated for it by Ren and Leskovec (2020). To further explore the potential of query answering in a different domain we use Hetionet (Himmelstein et al., 2017), a biomedical KG containing $\sim$2M edges between genes, diseases, and drugs, among other types. For Hetionet, we extract complex queries of different types, and then for both datasets we run the three steps above to collect preference data. Furthermore, in order to validate the alignment between preference sets derived from embeddings and human notions of similarity, we manually generate clusters of answers to queries for a subset of 140 queries in the FB15k237 dataset, and then measure the accuracy of clusters from embeddings with respect to the manually created ones, which results in an agreement of 79% (see Appendix˜B for details).
We consider 14 types of first order logic queries of varying complexity, including negation. We present statistics of the final datasets and splits in Table 1. For training NQR, we only use 1-hop, link prediction queries. For the validation and test sets we use the full set of 14 types of complex queries. More detailed statistics can be found in Appendix  C.